Allocating Joint Costs Using the Constant Gross Margin Method

A company manufactures three products, L-Ten, Triol, and Pioze, from a joint process. Each production run costs $12,900. None of the products can be sold at split-off, but must be processed further. Information on one batch of the three products is as follows:



Product
Gallons Further Processing
Cost per Gallon
Eventual Market
Price per Gallon
L-Ten 3,500 $0.50    $ 2.00   
Triol 4,000 1.00       5.00   
Pioze 2,500 1.50       6.00   

Required:

1. Calculate the total revenue, total costs, and total gross profit the company will earn on the sale of L-Ten, Triol, and Pioze.

Total Revenue $
Total Costs $
Total Gross Profit $

2. Allocate the joint cost to L-Ten, Triol, and Pioze using the constant gross margin percentage method. Round the gross margin percentage to four decimal places and round all other computations to the nearest dollar.

Joint Cost
Product Allocation
L-Ten $
Triol
Pioze
Total $

(Note: The joint cost allocation does not equal $12,900 due to rounding.)

3. What if it cost $2 to process each gallon of Triol beyond the split-off point? How would that affect the allocation of joint cost to these three products? Round the gross margin percentage to four decimal places and round all other computations to the nearest dollar.

Joint Cost
Product Allocation
L-Ten $
Triol
Pioze
Total $

(Note: The joint cost allocation does not equal $12,900 due to rounding.)
Answer #1
Please give positive ratings so I can keep answering. Thanks!
Answer 1
Product Gallons Further costs Further costs
L Ten          3,500.00                0.50        1,750.00
Triol          4,000.00                1.00        4,000.00
Pioze          2,500.00                1.50        3,750.00
Joint costs      12,900.00
Total costs      22,400.00
Product Gallons Market Price Market Value
L Ten          3,500.00                2.00        7,000.00
Triol          4,000.00                5.00      20,000.00
Pioze          2,500.00                6.00      15,000.00
Total Revenue      42,000.00 A
Total Revenue      42,000.00
Total costs      22,400.00
Total gross profit      19,600.00 B
Answer 2
Gross margin percentage 46.6667% C=B/A
(Total gross profit/ Total Revenue)
D E=D*C F G=D-E-F
Product Market Value Gross margin Further costs Joint costs allocated
L Ten          7,000.00         3,267.00        1,750.00     1,983.00
Triol        20,000.00         9,333.00        4,000.00     6,667.00
Pioze        15,000.00         7,000.00        3,750.00     4,250.00
Total 12,900.00
Answer 3
Product Gallons Further costs Further costs
L Ten          3,500.00                0.50        1,750.00
Triol          4,000.00                2.00        8,000.00
Pioze          2,500.00                1.50        3,750.00
Joint costs      12,900.00
Total costs      26,400.00 H
Total Revenue      42,000.00 See A
Total costs      26,400.00 See H
Total gross profit      15,600.00 I=A-H
Gross margin percentage 37.1429% J=I/A
(Total gross profit/ Total Revenue)
D K=D*J L M=D-K-L
Product Market Value Gross margin Further costs Joint costs allocated
L Ten          7,000.00         2,600.00        1,750.00     2,650.00
Triol        20,000.00         7,429.00        8,000.00     4,571.00
Pioze        15,000.00         5,571.00        3,750.00     5,679.00
Total 12,900.00